# Gradient-Regularized Out-of-Distribution Detection

### v1

@@ Abstract
One of the challenges for neural networks in real-life applications is the overconfident errors these models make when the data is not from the original training distribution. Addressing this issue is known as Out-of-Distribution (OOD) detection. Many state-of-the-art OOD methods employ an auxiliary dataset as a surrogate for OOD data during training to achieve improved performance. However, these methods fail to fully exploit the local information embedded in the auxiliary dataset. In this work, we propose the idea of leveraging the information embedded in the gradient of the loss function during training to enable the network to not only learn a desired OOD score for each sample but also to exhibit similar behavior in a local neighborhood around each sample. We also develop a novel energy-based sampling method to allow the network to be exposed to more informative OOD samples during the training phase. This is especially important when the auxiliary dataset is large. We demonstrate the effectiveness of our method through extensive experiments on several OOD benchmarks, improving the existing state-of-the-art FPR95 by $4\%$ on our ImageNet experiment. We further provide a theoretical analysis through the lens of certified robustness and Lipschitz analysis to showcase the theoretical foundation of our work. We will publicly release our code after the review process.
@@
**footnotetext: Equal contribution
Under review for the 18th European Conference on Computer Vision (ECCV) 2024
@@ 1 Introduction
Neural networks are increasingly being utilized across a wide range of applications and fields, achieving unprecedented levels of performance and surpassing traditional state-of-the-art approaches. However, concerns about robustness and safety, coupled with significant challenges in verifying robustness and ensuring safe performance impede their use in more sensitive applications. One concern that arises when deep models are deployed in the real world is their tendency to produce over-confident predictions upon encountering unfamiliar samples that are far distant from the space in which they were trained [39]. As a result, the field of Out-of-Distribution (OOD) detection has emerged to study and address this phenomenon.
@@
Many works have been presented in recent years, trying different methods and ideas to regulate the network’s output to OOD samples in terms of softmax probability [17, 18] or energy metrics [30, 35]. One family of methods, known as post-hoc approaches [47, 1, 28, 17, 31], operate on a given trained network and employ certain statistics, such as the confidence level of the network, the number of active neurons, the contribution of important pathways in the network, etc., to perform OOD detection without using any samples from an auxiliary distribution. Other methods [6, 30, 35, 22, 18] use OOD data for training or finetuning a network. The latter family usually tends to achieve better performances in comparison to post-hoc methods, as their access to the auxiliary dataset provides them with additional information regarding the outlier distribution.
@@ 2.2 OOD Detection with Auxiliary Outlier Dataset
When an auxiliary dataset is available, this group of methods [10, 17, 30] use this data to train the model to improve the network’s robustness against OOD data. OE [17] and Energy [30] propose their respective loss functions to be used during the training of the network. Recently, [7] improved [30] by proposing balanced Energy loss that balances the number of samples of the auxiliary OOD data across classes. Furthermore, [42] increases the expected Frobenius Jacobian norm difference between ID and OOD. This is in sharp contrast with our intuition as we aim to decrease the norm of the score function in all regions. Another line of work [11, 24, 14, 37, 36] focuses on generating artificial samples that resemble the real OOD data, using various generative models such as GANs. Most recently OpenMix [55] was proposed as a Misclassification detection method, teaching the model to reject pseudo-samples generated from the available outlier samples.
@@ 3.1 Notation
The In-Distribution (ID), Out-of-Distribution (OOD) and Auxiliary distribution over $\mathcal{X}$ are denoted by $\mathcal{D}_{\text{in}}$, $\mathcal{D}_{\text{out}}$ and $\mathcal{D}_{\text{aux}}$, respectively. The corresponding datasets are assumed to be sampled i.i.d from their respective distributions. We denote the indicator function of an event $e\leq 0$ as $\mathbb{I}_{e\leq 0}$, i.e., if $e\leq 0$ then $\mathbb{I}_{e\leq 0}=1$, and otherwise, $\mathbb{I}_{e\leq 0}=0$. We denote $\mathcal{B}(x,\varepsilon)=\{y|\|y-x\|\leq\varepsilon\}$. Finally, a function $f$ is said to be locally Lipschitz continuous on $\mathcal{X}$ if there exists a positive constant $L$ such that
$\|f(x)-f(y)\|\leq L\|x-y\|,\qquad\forall x,y\in\mathcal{X}.$
The smallest such constant is called the local Lipschitz constant of $f$.
@@ 4.2 OOD Sampling
One concern faced by OOD training schemes is that the OOD data may outnumber the ID data by orders of magnitude. Having access to such a large auxiliary dataset begs the question of which samples should the network see during the training phase. This creates two possible scenarios. Either the training algorithm exposes the model to an equal number of both ID and OOD data points, which means that the model cannot fully utilize the OOD dataset as most of the OOD dataset remains unseen. Alternatively, the algorithm may provide disproportionately more OOD samples, leading to bias in the model due to the imbalance in the samples [49].
These observations hint towards employing a more informative sampling strategy [27, 6] that utilizes as much of the dataset as possible while avoiding the pitfalls caused by imbalanced datasets. Whilst the first option that comes to mind is to perform a greedy sampling based on the score function $S$, it has been observed [22] that such greedy samples could bias the model to some specific regions of the image space, resulting in a model unable to generalize well.
To address these issues, we present Algorithm 1, which uses clustering (to discourage greedy behavior) alongside energy-based scoring to sample more informative images from the dataset. Next, we explain the sampling algorithm in more detail.
Clustering: Given samples $\{x_{i}\}_{i=1}^{n_{\text{OOD}}}$, we first calculate the features $z_{i}=h(x_{i})$ and the scores $s_{i}=-\text{LSE}(f(x_{i}))$. The next step is to perform clustering in the feature space. To perform clustering, we use K-Means [32] with a fixed number of clusters. As K-Means assigns clusters based on Euclidean distances, we normalize the features to $\hat{z}_{i}=\frac{z_{i}}{\|z_{i}\|_{2}}$ to avoid skewed clusters towards samples with disproportionate feature values.
@@ 5 Theoretical Analysis
where $L_{S}>0$. Suppose $x\sim\mathcal{D}_{\text{in}}$ is correctly classified by the score function as ID, i.e., $S(x)<\gamma$. If we want $x^{\prime}$ to be also labeled as ID, it is sufficient to satisfy
$S(x)+L_{S}\|x^{\prime}-x\|\leq\gamma,$
which in turn gives the following certificate
@@
Furthermore, for piecewise-linear networks like ReLU and LeakyReLU networks, the local Lipschitz constant is equal to the norm of the gradient. More specifically, for every point $x$***We discard the negligible case of points lying on the boundary of several piecewise-linear spaces., there exists an $\hat{\varepsilon}>0$ such that the network remains linear in this region, i.e.,
@@
Post-hoc Methods and Gradient Regularization: So far, we have discussed the idea of gradient regularization and provided a certified radius for which the model is robust against OOD misdetection using the Lipschitz constant of the network. This radius (4) is proportional to the inverse of the Lipschitz constant of the score function, and by regularizing the gradient we aim to reduce the Lipschitz constant to increase this radius. Here, we note that many of the state-of-the-art OOD detection methods can also be analyzed in this framework. For example, DICE [47] and LINe [1] use activation pruning and sparsification to remove noisy neurons. It has been shown [21] that for a given matrix, removing a row or column of it results in a reduction of the Lipschitz constant. As activation pruning is essentially a form of masking the weights of the network, methods utilizing activation pruning reduce the upper bound on the Lipschitz constant. By doing so, they acquire a network with a smaller Lipschitz upper bound that provides a larger certified radius.
@@ 6.1 Setup
Datasets. For CIFAR experiments, we use CIFAR-10 and CIFAR-100 as ID datasets. We evaluate the models on six different OOD datasets including Textures [8], SVHN [38], Places365 [54], LSUN-cropped, and LSUN-resized [52]. Following [47], at test time, we use 10000 randomly selected samples from each of these 6 datasets. Following [55], instead of the recently retracted TinyImages, we use the 300K RandomImages [18] as the auxiliary outlier dataset.
@@
For the ImageNet experiment we only perform finetuning on a pre-trained DenseNet-121 [19] due to limited resources. See the Supplementary Material for more details.
@@
Sampling Details: Following the explanations in Section 4.2, we cluster the auxiliary samples into $n_{\text{ID}}$ clusters and choose the lowest and highest scoring samples of each cluster, i.e., in each iteration, we present $2n_{\text{ID}}$ auxiliary samples for the training of that iteration. For the CIFAR experiments, we use the whole auxiliary dataset in each epoch for clustering. In other words, at each epoch, the 300K images are split into mini-batches, and clustering is performed on each mini-batch at each iteration so that no auxiliary sample is missed. This means the clustering is performed on each mini-batch and not the full auxiliary dataset, as doing so would be computationally prohibitive. For the ImageNet experiment, in each epoch, we sample 600K randomly selected images from the remaining (990) classes as the auxiliary dataset and perform the clustering on each mini-batch.
@@ 6.2 Results
Evaluation on CIFAR-10: We first evaluate the performance of our method on the CIFAR-10 dataset. Following the literature [1], Table 1 shows the average results over the six commonly used OOD datasets. The detailed results on individual OOD datasets can be found in the Supplementary Material. As the results on CIFAR-10 are saturated, in Table 1 we focus on the effectiveness of gradient regularization and report the rest of the methods in the Supplementary Material. It can be seen that GReg outperforms all the methods in both of the reported metrics. Furthermore, on CIFAR-10, the use of local information through gradient regularization, allows GReg to decrease the average FPR95 of Energy[30] by 3.9%.
Evaluation on CIFAR-100: For the results of this part, we use the same experimental setup and only change the ID dataset to CIFAR-100. The average results of this experiment set are presented in Table 2 and the individual results can be found in the Supplementary Material. In this setup, GReg provides competitive results on the ResNet and WRN architectures with state-of-the-art methods that do not employ sampling. Moreover, coupled with sampling, GReg+ outperforms all current state-of-the-art methods on almost all metrics of the different architectures. On average, GReg+ outperforms the FPR95 metric of DOS [22] by 1.8%, which showcases the effectiveness of energy-based sampling. The detailed results on each OOD dataset can be found in the Supplementary Material.
@@ 6.3 Ablation Study
Gradient Regularization: Due to the generality of the idea of gradient regularization, our method can be used in conjunction with many of the existing methods that require training. These methods include but are not limited to, Energy [30], OE [18], POEM [35], etc. In order to leverage the gradient regularization idea, one needs to define a suitable loss function based on the norm of the gradient of the score being used by the method. Once the gradient loss is designed, adding this term to the total loss function promotes a smoother score manifold learned by the neural network.
To show the effectiveness of gradient regularization on different OOD training schemes, we choose three of the most well-known OOD detection methods, Energy, OE, and POEM, and train a DenseNet with and without the gradient regularization loss on both of the CIFAR benchmarks. Table 4 shows the overall improvement of all these methods with gradient regularization. This improvement, in both the FPR95 and AUROC, supports our claim that regularizing the gradient improves the OOD detection performance and that coupling GReg with other methods could improve their performance.
Sampling Methods: To study the effect of various sampling methods, we compare POEM and DOS to our method with and without gradient regularization. We denote GReg+ without gradient regularization as Energy-only-clustering. Table 6 compares different methods that utilize some form of sampling and shows the impact of energy-based clustering. On top of that, the comparison between Energy-only-clustering and GReg+ further encourages the use of gradient regularization.
ID Accuracy: To study the effect of gradient regularization on accuracy, we compare the ID accuracy of our methods to the most relevant state-of-the-art on WRN and DenseNet architectures to ensure that our method does not have a large negative impact on the ID accuracy. Table 6 shows that gradient regularization (GReg) improves the ID accuracy of energy on both architectures and GReg+ also has superior ID accuracy in comparison with DOS.

### v3

@@ Abstract
One of the challenges for neural networks in real-life applications is the overconfident errors these models make when the data is not from the original training distribution. Addressing this issue is known as Out-of-Distribution (OOD) detection. Many state-of-the-art OOD methods employ an auxiliary dataset as a surrogate for OOD data during training to achieve improved performance. However, these methods fail to fully exploit the local information embedded in the auxiliary dataset. In this work, we propose the idea of leveraging the information embedded in the gradient of the loss function during training to enable the network to not only learn a desired OOD score for each sample but also to exhibit similar behavior in a local neighborhood around each sample. We also develop a novel energy-based sampling method to allow the network to be exposed to more informative OOD samples during the training phase. This is especially important when the auxiliary dataset is large. We demonstrate the effectiveness of our method through extensive experiments on several OOD benchmarks, improving the existing state-of-the-art FPR95 by $4\%$ on our ImageNet experiment. We further provide a theoretical analysis through the lens of certified robustness and Lipschitz analysis to showcase the theoretical foundation of our work. Our code is available at https://github.com/o4lc/Greg-OOD.
@@
**footnotetext: Co-first authors, equally contributed to the work
@@ 1 Introduction
Neural networks are increasingly being utilized across a wide range of applications and fields, achieving unprecedented performance levels and surpassing traditional state-of-the-art approaches. However, concerns about robustness and safety, coupled with significant challenges in verifying robustness and ensuring safe performance impede their use in more sensitive applications [32, 2]. One concern that arises when deep models are deployed in the real world is their tendency to produce over-confident predictions upon encountering unfamiliar samples that are distant from the space in which they were trained [47, 51]. As a result, the field of Out-of-Distribution (OOD) detection has emerged to study and address this phenomenon.
@@
Many works have been presented in recent years, trying different methods and ideas to regulate the network’s output to OOD samples in terms of softmax probability [22, 76, 48] or energy metrics [36]. One family of methods, known as post-hoc approaches [53, 41, 66, 72, 1, 49, 60, 34, 37], operate on a given trained network and employ certain statistics, such as the confidence level of the network, the number of active neurons, the contribution of important pathways in the network, etc., to perform OOD detection without using any samples from an auxiliary distribution. Other methods [7, 36, 43, 27, 23] use OOD data for training or finetuning a network. The latter family usually tends to achieve better performances in comparison to post-hoc methods, as their access to the auxiliary dataset provides them with additional information regarding the outlier distribution. OOD detection is closely related to anomaly [54, 69] and novelty detection [39, 10], each with slightly different problem formulations. Our work focuses on the general OOD detection problem.
@@ 2.2 OOD Detection with Auxiliary Outlier Dataset
When an auxiliary dataset is available, this group of methods [80, 13, 22, 14, 63, 36, 71, 68, 74] use this data to train the model to improve the network’s robustness against OOD data. OE [22] and Energy [36, 8] propose their respective loss functions to be used during the training of the network. Recently, [9] improved [36] by proposing balanced Energy loss that balances the number of samples of the auxiliary OOD data across classes. Furthermore, [52] increases the expected Frobenius Jacobian norm difference between ID and OOD. This is in sharp contrast with our intuition as we aim to decrease the norm of the score function in all regions. Another line of work [15, 29, 19, 45, 44] focuses on generating artificial samples that resemble the real OOD data, using various generative models such as GANs. Most recently, OpenMix [79] was proposed as a misclassification detection method, teaching the model to reject pseudo-samples generated from the available outlier samples.
@@ 3.1 Notation
The In-Distribution (ID), Out-of-Distribution (OOD) and Auxiliary distribution over $\mathcal{X}$ are denoted by $\mathcal{D}_{\text{in}}$, $\mathcal{D}_{\text{out}}$ and $\mathcal{D}_{\text{aux}}$, respectively. The corresponding datasets are assumed to be sampled i.i.d from their respective distributions. We denote the indicator function of an event $e\leq 0$ as $\mathbb{I}_{e\leq 0}$, i.e., if $e\leq 0$ then $\mathbb{I}_{e\leq 0}=1$, and otherwise, $\mathbb{I}_{e\leq 0}=0$. We denote $\mathcal{B}(x,\varepsilon)=\{y|\|y-x\|\leq\varepsilon\}$. Finally, a function $f$ is said to be locally Lipschitz continuous on $\mathcal{X}$ if there exists a positive constant $L$ such that $\|f(x)-f(y)\|\leq L\|x-y\|\ \forall x,y\in\mathcal{X}.$ The smallest such constant is called the local Lipschitz constant of $f$.
@@ 4.2 OOD Sampling
One concern faced by OOD training schemes is that the OOD data may outnumber the ID data by orders of magnitude. Having access to such a large auxiliary dataset begs the question of which samples should the network see during the training phase. This creates two possible scenarios. Either the training algorithm exposes the model to an equal number of both ID and OOD data points, which means that the model cannot fully utilize the OOD dataset as most of the OOD dataset remains unseen. Alternatively, the algorithm may provide disproportionately more OOD samples, leading to bias in the model due to the imbalance in the samples [62]. These observations hint towards employing a more informative sampling strategy [33, 7] that utilizes as much of the dataset as possible while avoiding the pitfalls caused by imbalanced datasets. Whilst the first option that comes to mind is to perform a greedy sampling based on the score function $S$, it has been observed [27] that such greedy samples could bias the model to some specific regions of the image space, resulting in a model unable to generalize well.
To address these issues, we present Algorithm 1, which uses clustering (to discourage greedy behavior) alongside energy-based scoring to sample more informative images. Next, we explain the sampling algorithm in more detail.
Clustering: Given samples $\{x_{i}\}_{i=1}^{n_{\text{OOD}}}$, we first calculate the features $z_{i}=h(x_{i})$ and the scores $s_{i}=-\text{LSE}(f(x_{i}))$. The next step is to perform clustering in the feature space. To perform clustering, we use K-Means with a fixed number of clusters. As K-Means assigns clusters based on Euclidean distances, we normalize the features to $\hat{z}_{i}=\frac{z_{i}}{\|z_{i}\|_{2}}$ to avoid skewed clusters towards samples with disproportionate feature values.
@@ 5 Theoretical Analysis
where $L_{S}>0$. Suppose $x\sim\mathcal{D}_{\text{in}}$ is correctly classified by the score function as ID, i.e., $S(x)<\gamma$. If we want $x^{\prime}$ to be also labeled as ID, it is sufficient to satisfy $S(x)+L_{S}\|x^{\prime}-x\|\leq\gamma,$ which in turn gives the following certificate
@@
Furthermore, for piecewise-linear networks like ReLU and LeakyReLU networks, the local Lipschitz constant is equal to the norm of the gradient. More specifically, for every point $x$*** We discard the negligible case of points on the boundaries of multiple piecewise-linear spaces., there exists an $\hat{\varepsilon}>0$ such that the network remains linear in this region, i.e.,
@@
Post-hoc Methods and Gradient Regularization: So far, we have provided a certified robustness radius against OOD misdetection using the Lipschitz constant of the network. This radius (4) is proportional to the inverse of the Lipschitz constant of the score function, and by regularizing the gradient we aim to reduce the Lipschitz constant to increase this radius. Here, we note that many of the state-of-the-art OOD detection methods can also be analyzed in this framework. For example, DICE [59] and LINe [1] use activation pruning and sparsification to remove noisy neurons. It has been shown [26] that for a given matrix, removing a row or column of it results in a reduction of the Lipschitz constant. As activation pruning is essentially a form of masking the weights of the network, methods utilizing activation pruning reduce the upper bound on the Lipschitz constant, acquiring a network with a smaller Lipschitz upper bound.
@@ 6.1 Setup
Datasets. For CIFAR experiments, we use CIFAR-10 and CIFAR-100 as ID datasets. We evaluate the models on six different OOD datasets including Textures [11], SVHN [46], Places365 [78], LSUN-cropped, LSUN-resized [70], and iSUN [67]. Following [59], at test time, we use 10000 randomly selected samples from each of these 6 datasets. Following [79], instead of the recently retracted TinyImages, we use the 300K RandomImages [23] as the auxiliary outlier dataset.
@@
For the ImageNet experiment, we only perform finetuning on a pre-trained DenseNet-121. See the Supplementary Material for more details.
@@
Sampling Details: Following the explanations in Section 4.2, we cluster the auxiliary samples into $n_{\text{ID}}$ clusters and choose the lowest and highest scoring samples of each cluster, i.e., in each iteration, we present $2n_{\text{ID}}$ auxiliary samples for the training of that iteration. For the CIFAR experiments, we use the whole auxiliary dataset in each epoch for clustering. In other words, at each epoch, the 300K images are split into mini-batches, and clustering is performed on each mini-batch at each iteration so that no auxiliary sample is missed. For the ImageNet experiment, in each epoch, we sample 600K randomly selected images from the remaining (990) classes as the auxiliary dataset and perform the clustering on each mini-batch.
@@ 6.2 Results
Evaluation on CIFAR-10: We first evaluate the performance of our method on the CIFAR-10 dataset. Following the literature [1], Table 1 shows the average results over the six commonly used OOD datasets. The detailed results on individual OOD datasets can be found in the Supplementary Material. As the results on CIFAR-10 are saturated, in Table 1 we focus on the effectiveness of gradient regularization and report the rest of the methods in the Supplementary Material. It can be seen that GReg outperforms all the methods in both of the reported metrics. Specifically, GReg decreases the FPR95 of [36] by 3.9%.
Evaluation on CIFAR-100: For the results of this part, we use the same experimental setup and only change the ID dataset to CIFAR-100. The average results of this experiment set are presented in Table 2 and the individual results can be found in the Supplementary Material. In this setup, GReg provides competitive results on the ResNet and WRN architectures with state-of-the-art methods that do not employ sampling. Moreover, coupled with sampling, GReg+ outperforms all current state-of-the-art methods on almost all metrics of the different architectures. On average, GReg+ outperforms the FPR95 metric of DOS [27] by 1.8%, which showcases the effectiveness of energy-based sampling. Results on each OOD dataset can be found in the Supplementary Material.
@@ 6.3 Ablation Study
Gradient Regularization: Due to the generality of the idea of gradient regularization, our method can be used in conjunction with many existing methods. To leverage the gradient regularization idea, one needs to define a suitable loss function based on the norm of the gradient of the score, to promote a smoother score manifold. To show the effectiveness of GReg on different OOD training schemes, we choose three well-known OOD detection methods, Energy, OE, and POEM, and train a DenseNet with and without gradient regularization on both CIFAR benchmarks. Table 4 shows the overall improvement of all these methods with gradient regularization, in both the FPR95 and AUROC, supporting our claim that coupling GReg with other methods could improve their performance.
Sampling Methods: To study the effect of various sampling methods, we compare POEM and DOS to our method with and without gradient regularization. We denote GReg+ without gradient regularization as Energy-only-clustering. Table 6 compares different methods that utilize some form of sampling and shows the impact of energy-based clustering. On top of that, the comparison between Energy-only-clustering and GReg+ further encourages gradient regularization.
ID Accuracy: To study the effect of gradient regularization on accuracy, we compare the ID accuracy of our methods to the most relevant state-of-the-art on WRN and DenseNet architectures. Table 6 shows that gradient regularization (GReg) improves the ID accuracy of energy on both architectures and GReg+ also has superior ID accuracy in comparison with DOS.
@@ 7 Conclusions
Acknowledgements
Research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-23-2-0008. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.